OpenStudy (anonymous):

A motorcyclist is driving up a ramp that has a constant incline relative to level ground. After he rides 1000 feet on the ramp, he is 100 feet off the ground. What is his speed relative to the ground if he drives at 80 miles per hour up the ramp?

OpenStudy (amistre64):

using a right triangle, we sort of relationship can we construct to play with?

OpenStudy (amistre64):

*what sort of ...

OpenStudy (anonymous):

I think it has to do with the derivative

OpenStudy (amistre64):

it does, but we need to relate the information to something more equationy, like may pythag thrm

OpenStudy (amistre64):

spped relative to the ground, is that his horizontal speed? or vertical spped?

OpenStudy (anonymous):

I already solved for these values 

OpenStudy (anonymous):

The height is 100 btw (it got cut off in the drawing)

OpenStudy (amistre64):

it shows on my screen, might be a zoom thing on yours

OpenStudy (amistre64):

well speed is a linear factor just as length is ... so whatever direction is being requested has the same ratio 1000 80 ----- = ------- 100 vertical or the other leg for the horizontal

OpenStudy (amistre64):

or by trig: 80 cos(theta) or 80 sin(theta) are just as good

OpenStudy (anonymous):

I don't really understand

OpenStudy (amistre64):

 its just similar triangles

OpenStudy (anonymous):

I'm not sure that is what the question is asking. I am pretty sure it has to do with the derivative because that is what we learned in class.

OpenStudy (amistre64):

well it does have to do with the derivative, but it can be simplified to a rather basic algebra/trig problem

OpenStudy (anonymous):

Can you explain how to use the derivative to get the answer, so I can understand the question in terms with what we are learning in class?

OpenStudy (amistre64):

assuming speed relative to the ground is horizontal, or in the +x direction cos(a) = x/r take implicit derivative to get -sin(a) a' = (rx'-xr')/r^2 since the angle doesnt change, a'=0; and r' is given as 80, r=1000 0 = (1000x'-80x)/(1000)^2 you determined x is about 995 sooo 0 = (1000x'-80(995))/(1000)^2 solve for x'

OpenStudy (amistre64):

80 (995) = 1000 x' which of course is the same ratio of similar triangles: 1000/995 = 80/x'

OpenStudy (anonymous):

So x' would be 398/5?

OpenStudy (amistre64):

if relative to the ground is in the +y direction sin(a) = y/r same process

OpenStudy (anonymous):

I just realized something. At the beginning, wouldn't we have to convert 80 mph to ft/s?

OpenStudy (amistre64):

no

OpenStudy (anonymous):

Oh okay. Just making sure. Thank you for your help.

OpenStudy (amistre64):

the angle of the ramp is not changing, so all linear forces/measures define similar triangles ....

OpenStudy (amistre64):

good luck